OpenStudy (anonymous):

((This is probably a really easy problem but i need help pls)) What is the simplest form of 16^3/4? a.8 b.12 c.24 d.64

jimthompson5910 (jim_thompson5910):

16^(3/4) 16^(3*1/4) 16^(1/4*3) [ 16^(1/4) ] ^3 [ 4th root of 16 ] ^3 [ 2 ] ^3 8 Hopefully all that makes sense. If not, then let me know

jimthompson5910 (jim_thompson5910):

I'm using the ideas that x^(y*z) = (x^y)^z and x^(1/n) = nth root of x

OpenStudy (anonymous):

woah okay, take that a little slower omg

jimthompson5910 (jim_thompson5910):

sorry about that

jimthompson5910 (jim_thompson5910):

which step isn't making sense?

OpenStudy (anonymous):

16^(3/4) is equal to \[\sqrt[4]{16}^{3}\] because \[x ^{\frac{ n }{ m }} = (\sqrt[m]{x})^{n}\] 4th root of 16 is 2 because 2x2x2x2 = 16. So you end up with\[2^{3} = 8 \]

OpenStudy (anonymous):

what are you doing to simplify 3/4

jimthompson5910 (jim_thompson5910):

I turned 3/4 into 3*1/4 I wanted to pull out that 3 so to speak so I can be left with 1/4 then I'd use the rule that TripleC just posted

OpenStudy (anonymous):

@jim_thompson5910 Who needs to pull it out when you can do it in one big bang?

jimthompson5910 (jim_thompson5910):

that's true, you have a much simpler/shorter way to do it

OpenStudy (anonymous):

ok ok, so if it was another one like idk 24 ^5/6, you would make it, ^6 sqrt 24 ^5??

OpenStudy (anonymous):

\[(\sqrt[6]{24})^{5}\]

OpenStudy (anonymous):

That just needs bunging into a calculator or left alone as an evaluated form.

OpenStudy (anonymous):

So yeah you were right xD

OpenStudy (anonymous):

alright, this makes more sense haha

OpenStudy (anonymous):

@jim_thompson5910 can you help with another one really quick??

jimthompson5910 (jim_thompson5910):

sure, what do you need

OpenStudy (anonymous):

solution for x= sqrt -x+6

jimthompson5910 (jim_thompson5910):

so the equation is \[\large x = \sqrt{-x+6}\] right?

OpenStudy (anonymous):

yes (:

jimthompson5910 (jim_thompson5910):

ok great

jimthompson5910 (jim_thompson5910):

first you square both sides to get rid of that square root, then you get everything to one side, then factor like so \[\large x = \sqrt{-x+6}\] \[\large x^2 = (\sqrt{-x+6})^2\] \[\large x^2 = -x+6\] \[\large x^2 + x-6=0\] \[\large (x+3)(x-2)=0\] I'll let you finish. Make sure you check all of your possible answers.

OpenStudy (anonymous):

would you distribute, so it would be like x^2-6??

jimthompson5910 (jim_thompson5910):

no you use the zero product property so (x+3)(x-2)=0 turns into x+3 = 0 or x-2 = 0 then you solve each for x to get x = -3 or x = 2 I'll let you check these answers

OpenStudy (anonymous):

okay, so would the zero property change whether its positive or negative

jimthompson5910 (jim_thompson5910):

the zero product property is the idea that if A*B = 0 then A = 0 or B = 0

OpenStudy (anonymous):

okay because for the problem, it was 3 and -2, but the answer was -3 and 2

jimthompson5910 (jim_thompson5910):

well that's because when you solve x+3 = 0, you get x = -3 see how that works out?

OpenStudy (anonymous):

oh okay i got it(:

jimthompson5910 (jim_thompson5910):

I'm glad you do

OpenStudy (anonymous):

thank you so much ((:

jimthompson5910 (jim_thompson5910):

you're welcome